OpenStudy (anonymous):

Given the following linear equations: 2x+3y=10 4x-6y=-28 Solve for (x,y) if there is a solution

OpenStudy (watchmath):

Mutiply the first equation by 2 and add it to the second to get 8x=-8 x=-1 -2+3y=10 3y=12 y=4 (-1,4)

OpenStudy (anonymous):

its very simple

OpenStudy (anonymous):

Multiply the first equation by 2 and add it to the second equation to get 8x=-8 Therefore x=-1 3y=10-2x = 10-2(-1) = 10+2 = 12 Therefore y = 12 / 3 = 4 (-1,4)